One of two numbers 9,4is AM of two positive numbers and other is gmof these numbers.  These numbers are the roots of the equation 

Grade:12th pass

1 Answers

Vikas TU
4 years ago
Hiii 
A.M means arithmetic mean is given by 
(a+b)/2 = 9 
GM means geometric mean is given by 
sqrt (ab) = 4 
ab = 16 
a = 18 – b 
Put the value of a 
(18-b) b = 16 
b^2 + 16 -18b = 0 
It will be having 2 roots 9+aqrt(65)
and 9- sqrt(65)